# Harmonized Speculative Sampling

### v1

@@ Abstract
Speculative sampling has proven to be an effective solution to accelerate decoding from large language models, where the acceptance rate significantly determines the performance. Most previous works on improving the acceptance rate focus on aligned training and efficient decoding, implicitly paying less attention to the linkage of training and decoding. In this work, we first investigate the linkage of training and decoding for speculative sampling and then propose a solution named HArmonized Speculative Sampling (HASS). HASS improves the acceptance rate without extra inference overhead by harmonizing training and decoding on their objectives and contexts. Experiments on three LLaMA models demonstrate that HASS achieves 2.81x-3.65x wall-clock time speedup ratio averaging across three datasets, which is 8%-15% faster than EAGLE-2.
@@ 1 Introduction
Generative large language models (LLMs), such as GPT-4 (Achiam et al., 2023) and LLaMA (Touvron et al., 2023), achieve impressive capabilities on various tasks. However, it is challenging to efficiently decode from LLMs because of the auto-regressive decoding mechanism, limiting their usage to time-sensitive applications. Speculative sampling (Chen et al., 2023a; Leviathan et al., 2023) addresses this by utilizing more concurrency from extra resources. In particular, it employs an efficient draft model to generate draft tokens auto-regressively and then verifies them by the target LLM in parallel. According to the verification result, a subset of draft tokens maintaining the same distribution of the target LLM is accepted as the decoding output.
Leviathan et al. (2023) shows that the higher acceptance rate leads to more wall-clock time acceleration and less extra computation resources. Most of previous works on improving the acceptance rate fall into two categories: aligned training (Li et al., 2024a; Zhou et al., 2023; Liu et al., 2023) and efficient decoding (Li et al., 2024b; Miao et al., 2024; Sun et al., 2024; Chen et al., 2023b). The former aims to obtain well-aligned draft models that behave similarly to the target LLM before the decoding stage, while the latter tries to gain more concurrency efficiently during the decoding stage. However, these works usually consider the training and decoding separately, paying less attention to the linkage between them on their objectives and contexts.
@@
During the decoding stage, the objective of draft models is to propose tokens that the target LLM may give high probabilities (Li et al., 2024b; Miao et al., 2024; Sun et al., 2024). In this case, the draft model should spend more on the recall of desired tokens, and the order of these tokens can be appropriately weakened. On the other hand, most LLM applications perform nucleus sampling (Holtzman et al., 2019) or top-k sampling (Fan et al., 2018). For such decoding objectives, tokens with high probabilities have a more significant impact on the result. Consequently, with precious capacity to build efficient draft models, the training objective of draft models should be aware of above properties in the decoding stage. To the best of our knowledge, previous works on training draft models for speculative sampling are independent of these decoding properties.
To demonstrate the disharmony of training and decoding on contexts, we take EAGLE (Li et al., 2024a) as an example. EAGLE employs previous hidden states of the target LLM to generate subsequent draft tokens, which performs excellent acceptance rates with an efficient draft model. Note that other approaches utilizing hidden states of the target LLM, such as GLIDE (Du et al., 2024) and Clover (Xiao et al., 2024), are also suitable for the following analysis. As shown in Figure 2, the draft model always has access to the target LLM’s hidden states in previous timesteps during training. However, during the decoding stage, the draft model can not access the target LLM’s hidden states about unverified timesteps, leading to the context misalignment of training and decoding. One can view this issue as a kind of exposure bias (Bengio et al., 2015; Wang & Sennrich, 2020) at the feature level in speculative sampling.
@@
Figure 1: Speedup ratios of different methods on LLaMA2-Chat 7B/13B and LLaMA3-Instruct 8B with temperature T $\in\{0,1\}$, averaging over MT-bench, HumanEval, and GSM8K datasets.
@@
Figure 2: We use EAGLE (Li et al., 2024a) as an example to illustrate the context misalignment, where the speculation starts from timestep $t$. $f^{(l)}$ and $f^{(s)}$ represent hidden states from the target LLM and the draft model, respectively. When decoding draft token $x_{t+2}$, the input context is inconsistent between training and decoding.
@@
In this paper, we thus propose a solution named HArmonized Speculative Sampling (HASS) that improves the acceptance rate by harmonizing training and decoding on their objectives and contexts. Specifically, to make draft models aware of the decoding strategy, HASS extends the idea about ranking distillation (Tang & Wang, 2018) from the recommender system to speculative sampling, resulting in a distillation loss concerning most probable tokens w.r.t. the target distribution. To harmonize training and decoding on their contexts, HASS employs a context-aligned training strategy to address the context misalignment discussed before. These two strategies of HASS improve the acceptance rate without any inference overhead and are also efficient during training.
@@
We conduct experiments across dialogue, code generation, and mathematical reasoning using the MT-bench, HumanEval, and GSM8K datasets, respectively. LLaMA2-Chat 7/13B and LLaMA3-Instruct 8B are considered as target LLMs. Building with EAGLE-2 (Li et al., 2024b), HASS achieves 8%-12% acceptance length improvement over EAGLE-2, resulting in 2.81x-3.65x wall-clock time acceleration compared with the vanilla inference on a single NVIDIA H800 GPU.
@@ 2 Preliminary
Speculative sampling is an implementation of speculative execution (Kung & Robinson, 1981; Hennessy & Patterson, 2011) for reducing wall-clock time by utilizing more concurrency. Specifically, given the target LLM $\mathcal{M}^{(l)}$ intended for acceleration, speculative sampling employs a draft model $\mathcal{M}^{(s)}$ to speculatively and efficiently generate draft tokens. The standard approach (Leviathan et al., 2023; Chen et al., 2023a) breaks down the next step generation into three steps:
@@
  • $\tau$ tokens that maintain the target distribution are accepted by a modified rejection sampling schema based on the draft sequence and the distribution gap.
Leviathan et al. (2023) proves that the wall-clock time improvement ratio is proportional to $\tau$ and the arithmetic operation increment ratio is inversely proportional to $\tau$. Consequently, $\tau$, known as the acceptance length, is vital for the acceleration performance. Similar analysis is also suitable when utilizing multiple draft sequences (Miao et al., 2024; Li et al., 2024b; a; Sun et al., 2024).
Note that $\tau$ is highly related to the distribution gap between the target LLM and the draft model. With efficient decoding requirements, the draft model usually has limited capacity, leading to an unbridgeable distribution gap with the target LLM. Fortunately, during inference, the acceptance rate is mainly determined by the distribution alignment on desired tokens, i.e., tokens that the target LLM gives high probabilities. However, previous speculative sampling works mainly focus on the entire vocabulary set w.r.t. knowledge distillation from the target LLM (Li et al., 2024a; Zhou et al., 2023), which separates the training from the practical requirements of the decoding.
EAGLE (Li et al., 2024a) is a lightweight draft model design, as shown in Figure 2. During decoding, EAGLE utilizes the LM Head of the target LLM to generate draft tokens. Specifically, suppose that the speculation starts from timestep $t$, i.e., the first draft token is at timestep $t+1$. When generating draft token $x_{t+1}$, target LLM’s hidden state $f_{t-1}^{(l)}$ in second-to-top-layer is concatenated with the embedding $e_{t}$ to perform the input of the draft model. During training, EAGLE constructs a regression task between $f^{(l)}$s and the draft model’s predicted hidden states $f^{(s)}$s.
However, because of the auto-regressive decoding, the draft model only accesses the target LLM’s features at the start of the speculation. EAGLE employs produced features of draft models as input in the following steps. This context misalignment caused by the feature inaccuracy introduces error accumulation and impairs the performance of generating later draft tokens. EAGLE-2 (Li et al., 2024b) employs the same model design but works on dynamic drafting structures instead of a static tree structure during the decoding stage, still leaving the above issue.
@@ 3 Methodology
As described in Sec. 2, previous speculative sampling methods focus on one of the training and decoding, ignoring the linkage on their objectives and contexts. In this section, we introduce HArmonized Speculative Sampling (HASS) to address the disharmony issues above. In particular, HASS consists of two parts: harmonized objective distillation and harmonized context alignment.
@@ 3.1 Harmonized Objective Distillation
HASS pays more attention to desired tokens by leveraging the ranking distillation (Tang & Wang, 2018) idea from the recommender system. In particular, ranking distillation aims to learn a student model that gives higher ranks for the teacher model’s top ranking of items. In speculative sampling, the draft model and the target LLM act as the student and the teacher, respectively. Draft models with similar properties will perform at a higher acceptance rate in the decoding stage. Let the set of $K$ tokens with the highest probabilities from the target LLM’s probability distribution as $\hat{\Omega}\subset\Omega$, where $\Omega$ is the entire vocabulary. HASS considers the following Top-$K$ distillation loss:
@@
where $q$ and $p$ are the next token probability distributions of the target LLM and the draft model, respectively. The training objective above is designed to generate tokens that are friendly to the decoding strategies without extra computation overhead during inference. Note that, when building with EAGLE, the training stage can obtain $\hat{\Omega}$ from hidden states of the target LLM, i.e., the proposed loss also enjoys the efficient training cost as EAGLE.
@@ 3.2 Harmonized Context Alignment
Figure 3: Training with harmonized context alignment, where $q$ and $k$ refer to the query and key states in the transformer layer, respectively. Superscript $(l)$ denotes tensors from the target LLM, and superscript $(s_{j})$ denotes tensors from the $j$-th draft model forward. Note that during training $(s_{j})$ refers to call $j$ times draft model in a batch, while during inference $(s_{j})$ refers to $j$-th auto-regressive decoding.
HASS follows a context alignment schema that harmonizes training and decoding on their contexts. In particular, HASS divides the training procedure into $n$ steps so that the draft model can access contextual features that are consistent with the decoding stage, as described below:
  • The first step is the same as the training stage of EAGLE. Given timestep $t+1$, the draft model takes the target model’s hidden feature $f^{(l)}_{t}$ as input and produces the draft feature $f^{(s_{1})}_{t+1}$. Here, the attention mask is the original causal mask without modification.
  • The second step takes the first step’s features into account. For instance, in the self-attention mechanism of timestep $t+1$, $f^{(s_{1})}_{t}$ is used to calculate the current query, with keys and values from $f^{(l)}_{:t}\oplus f^{(s_{1})}_{t}$, where $\oplus$ refers to concatenate and $f^{(l)}_{:t}$ means features in timesteps less than $t$. We modify the attention mask to ensure that the previous feature seen by $f^{(s_{1})}_{i}$ is always $f^{(l)}_{i-1}$, as shown in the ‘HASS Training Step 2’ part of Figure 3.
  • For step $j\geq 3$, the feature from the last step $f^{(s_{j-1})}_{t}$ is used to calculate the query in timestep $t+1$, while $f^{(l)}_{:t-j+2}\oplus f^{(s_{1})}_{t-j+2}\oplus\ldots\oplus f^{(s_{j-1})}_{t}$ is used to generate keys and values in the self-attention mechanism.
It takes $n$ times the training overhead of EAGLE and keeps the same decoding overhead. We empirically demonstrate that the acceleration effect is converged with a small $n$ so that the training of HASS is cost-efficient.
@@ 4.1 Experimental Setup
Target LLMs. LLaMA2-Chat 7B/13B and LLaMA3-Instruct 8B.
@@
Metrics. HASS neither fine-tunes the target LLMs’ weights during training nor relaxes the acceptance conditions during decoding, making it a lossless acceleration method. Thus, the generation quality is promised with no need for evaluation. We use the following two metrics to assess the acceleration performance:
  • Speedup Ratio: The actual test speedup ratio w.r.t. vanilla auto-regressive decoding.
  • Acceptance Length $\tau$: The average number of tokens generated per drafting-verification cycle corresponds to the number of tokens accepted by the target model from the draft model.
Note that the speedup ratio is sensitive to the hardware because of the computing power, and the acceptance length is also slightly affected by the hardware because of the numerical error. Therefore, we perform all the inference processes on a single NVIDIA H800 GPU.
Comparisons. The vanilla auto-regressive decoding is taken as the baseline, which serves as the benchmark for speedup ratios (1.00x). We compare HASS with recent lossless speculative sampling methods, including PLD (Saxena, 2023), Lookahead (Fu et al., 2023), EAGLE (Li et al., 2024a), and EAGLE-2 (Li et al., 2024b).
Implementation. Our code is built based on EAGLE-2’s open source repository [^1] https://github.com/SafeAILab/EAGLE. Experiments on EAGLE and EAGLE-2 reuse draft model weights trained by Li et al. (2024a). For harmonized objective distillation, $K$ is set as 10, and the loss of harmonized objective distillation is added to EAGLE’s original loss with a coefficient of 1.0. For harmonized context alignment, the draft model is aligned for 3 steps during training. For dynamic tree structure, we set the total number of draft tokens to 60 for all experiments with a draft tree depth of 6. We keep other settings, such as the fixed training dataset and the optimizer, consistent with EAGLE-2.

### v2

# Learning Harmonized Representations for Speculative Sampling

@@ Abstract
Speculative sampling is a promising approach to accelerate the decoding stage for Large Language Models (LLMs). Recent advancements that leverage target LLM’s contextual information, such as hidden states and KV cache, have shown significant practical improvements. However, these approaches suffer from inconsistent context between training and decoding. We also observe another discrepancy between the training and decoding objectives in existing speculative sampling methods. In this work, we propose a solution named HArmonized Speculative Sampling (HASS) that learns harmonized representations to address these issues. HASS accelerates the decoding stage without adding inference overhead through harmonized objective distillation and harmonized context alignment. Experiments on four LLaMA models demonstrate that HASS achieves 2.81x-4.05x wall-clock time speedup ratio averaging across three datasets, surpassing EAGLE-2 by 8%-20%.
@@ 1 Introduction
Generative Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023) and LLaMA (Touvron et al., 2023), have demonstrated remarkable capabilities across a wide range of tasks. Nevertheless, efficiently decoding from these models poses a significant challenge due to the inherent auto-regressive decoding mechanism, which restricts their applicability in time-sensitive scenarios. Speculative sampling (Chen et al., 2023; Leviathan et al., 2023) offers a solution by leveraging additional resources to increase concurrency. Specifically, it employs an efficient draft model to generate draft tokens auto-regressively, which are then concurrently verified by the target LLM. Based on the verification results, a subset of draft tokens that preserves the same distribution as the target LLM is accepted as the final output.
Leviathan et al. (2023) show that the practical performance of speculative sampling is highly related to two factors: the decoding cost of the draft model and its alignment with the target LLM. To develop efficient draft models that are well-aligned with the target LLM, previous works propose to leverage the target LLM’s contextual information (Xiao et al., 2024; Li et al., 2024b; c; Du et al., 2024). For instance, EAGLE (Li et al., 2024b; c) employs previous hidden states of the target LLM as the draft model’s input features. However, these approaches introduce inconsistent context between training and decoding, as illustrated in Figure 2. During training, the draft model always has access to the target LLM’s hidden states in previous timesteps. However, during decoding, the draft model cannot access the target LLM’s hidden states for unverified timesteps, resulting in a context misalignment between training and decoding. This issue can be viewed as a form of exposure bias (Bengio et al., 2015; Wang & Sennrich, 2020) at the feature level in speculative sampling.
@@
Another discrepancy is also observed between the objectives of the training and decoding stages. During the decoding stage, the objective of the draft model is to propose tokens that the target LLM is likely to assign high probabilities to (Li et al., 2024c; Miao et al., 2024; Sun et al., 2024). In this scenario, the draft model should focus more on recalling the desired tokens, while the specific order of these tokens can be somewhat de-emphasized. Moreover, most LLM applications perform nucleus sampling (Holtzman et al., 2019) or top-k sampling (Fan et al., 2018). For these decoding objectives, tokens with high probabilities play a more significant role in determining the output. Therefore, to develop efficient draft models, their training objectives should consider these properties encountered in the decoding stage. To the best of our knowledge, previous works on training draft models for speculative sampling have largely overlooked these decoding considerations.
@@
Figure 1: Speedup ratios of different methods on LLaMA2-Chat 7B/13B and LLaMA3-Instruct 8/70B with temperature T $\in\{0,1\}$, averaging over MT-bench, HumanEval, and GSM8K datasets.
@@
Figure 2: We use EAGLE (Li et al., 2024b) as an example to illustrate the context misalignment, where the speculation starts from timestep $t$. $f^{(l)}$ and $f^{(s)}$ represent hidden states from the target LLM and the draft model, respectively. When decoding draft token $x_{t+2}$, the input context is inconsistent between training and decoding.
@@
In this paper, we introduce HArmonized Speculative Sampling (HASS), a novel approach designed to address the aforementioned issues by learning harmonized representations. Specifically, to make draft models aware of the decoding strategy, HASS extends the idea of ranking distillation (Tang & Wang, 2018) from the recommender system to speculative sampling, resulting in a distillation loss focused on the most probable tokens within the target distribution. To mitigate the previously discussed context misalignment between training and decoding, HASS employs a context-aligned training strategy. Together, these two strategies of HASS improve the acceleration performance without any inference overhead and maintain training efficiency.
@@
We conduct experiments across dialogue, code generation, and mathematical reasoning using the MT-bench, HumanEval, and GSM8K datasets, respectively. Building with EAGLE-2 (Li et al., 2024c), HASS achieves 8%-16% acceptance length improvement over it on LLaMA2-Chat 7/13B and LLaMA3-Instruct 8/70B, resulting in 2.81x-4.05x wall-clock time acceleration compared with the vanilla inference on NVIDIA H800 GPU.
@@ 2 Preliminary
Speculative sampling leverages the concept of speculative execution (Kung & Robinson, 1981; Hennessy & Patterson, 2011) to reduce wall-clock time from more concurrency. Specifically, given the target LLM $\mathcal{M}^{(l)}$ that is the focus of acceleration, speculative sampling employs a draft model $\mathcal{M}^{(s)}$ to speculatively and efficiently generate draft tokens. The conventional approach (Leviathan et al., 2023; Chen et al., 2023) decomposes the next step generation into three steps:
@@
  • $\tau$ tokens that retain the target distribution are accepted by a modified rejection sampling schema based on the draft sequence and the distribution gap.
Leviathan et al. (2023) demonstrate that the wall-clock time improvement ratio is directly proportional to $\tau$, while the arithmetic operation increment ratio is inversely proportional to $\tau$. Consequently, $\tau$, also known as the acceptance length, plays a crucial role in determining the performance of acceleration. This analysis also applies when using multiple draft sequences (Miao et al., 2024; Li et al., 2024c; b; Sun et al., 2024). Note that $\tau$ is closely related to the distribution gap between the target LLM and the draft model. With efficient decoding requirements, the draft model typically has limited capacity, resulting in a significant distribution gap compared to the target LLM. Fortunately, during inference, the acceptance rate is primarily influenced by the alignment of distributions on the desired tokens, i.e., the tokens to which the target LLM assigns high probabilities. However, previous speculative sampling works mainly focus on the entire vocabulary set w.r.t. knowledge distillation from the target LLM (Li et al., 2024b; Zhou et al., 2023), thereby disconnecting the training process from the practical decoding requirements.
EAGLE (Li et al., 2024b) is a lightweight draft model design, as shown in Figure 2. During decoding, it utilizes the LM Head of the target LLM to generate draft tokens. Specifically, we assume that the speculation starts from timestep $t$, meaning the first draft token is at timestep $t+1$. To generate the draft token $x_{t+1}$, the target LLM’s hidden state $f_{t-1}^{(l)}$ in the second-to-top layer is concatenated with the embedding $e_{t}$ to perform the input of the draft model. During training, EAGLE constructs a regression task between $f^{(l)}$s and the predicted hidden states $f^{(s)}$s of the draft model. However, due to the auto-regressive decoding, the draft model only accesses the target LLM’s features at the beginning of the speculation. It uses the features produced by the draft models as input for subsequent steps. This context misalignment, stemming from feature inaccuracies, leads to error accumulation and hinders the performance of generating later draft tokens (Li et al., 2024b; Du et al., 2024). EAGLE-2 (Li et al., 2024c) employs the same model design but works on dynamic drafting structures instead of a static tree structure during the decoding stage, yet the aforementioned issue remains unresolved.
@@ 3 Methodology
As outlined before, previous speculative sampling methods suffer from disharmonies between training and decoding. This section introduces HArmonized Speculative Sampling (HASS) to tackle objective misalignment and context inconsistency through harmonized objective distillation and harmonized context alignment, respectively, as described below.
@@ 3.1 Harmonized Objective Distillation
HASS prioritizes the most decoding-desired tokens by leveraging the ranking distillation (Tang & Wang, 2018) idea from the recommender system. Specifically, ranking distillation aims to train a student model to assign higher ranks to the items that are top-ranked by the teacher model. In the context of speculative sampling, the draft model and the target LLM serve as the student and the teacher, respectively. Draft models with similar properties will perform at a higher acceptance rate in the decoding stage. Consider the set of K tokens with the highest probabilities from the target LLM’s probability distribution as $\hat{\Omega}\subset\Omega$, where $\Omega$ represents the entire vocabulary. HASS considers the following Top-K distillation loss:
@@
where $q$ and $p$ are the next token probability distributions of the target LLM and the draft model, respectively. Note that, when integrated with EAGLE, the training stage can obtain $\hat{\Omega}$ from hidden states of the target LLM. This implies that the proposed loss function benefits from the same efficient training cost as EAGLE. We evaluate the proposed Top-K distillation loss against six alternative losses, such as BiLD (Li et al., 2024a) and Recall@k Surrogate loss (Patel et al., 2022), through ablation studies.
@@ 3.2 Harmonized Context Alignment
Figure 3: Training with harmonized context alignment, where $q$ and $k$ refer to the query and key states in the transformer layer, respectively. Superscript $(l)$ denotes tensors from the target LLM, and superscript $(s_{j})$ denotes tensors from the $j$-th draft model forward. Note that during training $(s_{j})$ refers to calling $j$ times draft model in a batch, while during inference $(s_{j})$ refers to $j$-th auto-regressive decoding.
HASS follows a context alignment schema that aligns training and decoding on their contexts. In particular, HASS divides the training procedure into $n$ steps, enabling the draft model to utilize contextual features consistent with those in the decoding stage. The procedure is outlined as follows:
  • The first step mirrors the training stage of EAGLE. At timestep $t+1$, the draft model takes the target model’s feature $f^{(l)}_{t}$ as input and produces the draft feature $f^{(s_{1})}_{t+1}$. In this step, the attention mask remains as the original causal mask without any modification.
  • In the second step, features from the first step are incorporated. For instance, in the self-attention mechanism at timestep $t+1$, $f^{(s_{1})}_{t}$ is used to calculate the current query. Keys and values are derived from $f^{(l)}_{:t}\oplus f^{(s_{1})}_{t}$, where $\oplus$ denotes concatenation and $f^{(l)}_{:t}$ includes features from timesteps earlier than $t$. The attention mask is adjusted to ensure that the previous feature seen by $f^{(s_{1})}_{i}$ is always $f^{(l)}_{i-1}$, as shown in the ‘HASS Training Step 2’ part of Figure 3.
  • For step $j\geq 3$, the feature from the previous step $f^{(s_{j-1})}_{t}$ is utilized to calculate the query at timestep $t+1$, while keys and values are generated by $f^{(l)}_{:t-j+2}\oplus f^{(s_{1})}_{t-j+2}\oplus\ldots\oplus f^{(s_{j-1})}_{t}$.
It takes $n$ times the training overhead of EAGLE while maintaining the same decoding overhead. We empirically demonstrate that the acceleration effect converges with a small $n$ so that the training of HASS is cost-efficient.
@@ 4.1 Experimental Setup
Target LLMs. LLaMA2-Chat 7B/13B and LLaMA3-Instruct 8/70B.
@@
Metrics. HASS neither fine-tunes the target LLMs’ weights during training nor relaxes the acceptance conditions during decoding, making it a lossless acceleration method. Thus, the generation quality is promised with no need for evaluation. We use the following two metrics to measure the acceleration performance:
  • Speedup Ratio: The actual test speedup ratio relative to vanilla auto-regressive decoding.
  • Acceptance Length $\tau$: The average number of tokens generated per drafting-verification cycle, indicating the number of tokens accepted by the target LLM from the draft model.
Note that the speedup ratio is sensitive to the hardware due to variations in computing power, and the acceptance length may be slightly affected by hardware due to numerical errors. Therefore, all inference processes are conducted on NVIDIA H800 GPU.
Comparisons. The vanilla auto-regressive decoding is taken as the baseline, which serves as the benchmark for speedup ratios (1.00x). We compare HASS with recent lossless speculative sampling methods, including PLD (Saxena, 2023), Lookahead (Fu et al., 2023), EAGLE (Li et al., 2024b), and EAGLE-2 (Li et al., 2024c).
Implementation. Our code is built based on EAGLE-2’s open-source repository [^1] https://github.com/SafeAILab/EAGLE. Experiments on EAGLE and EAGLE-2 reuse draft model weights trained by Li et al. (2024b). For harmonized objective distillation, K is set as 10, and the loss of harmonized objective distillation is added to EAGLE’s original loss with a coefficient of $w=1.0$. For harmonized context alignment, the draft model is aligned for 3 steps during training. For dynamic tree structure, we set the total number of draft tokens to 60 for all experiments with a draft tree depth of 6. We keep other settings, such as the fixed training dataset and the optimizer, consistent with EAGLE-2.